# Fast nonparametric inference of network backbones for graph sparsification

### v1

@@ III.1 Synthetic Backbone Reconstruction
Figure 2: Reconstruction of planted backbone structure in synthetic network data. (a) Network Jaccard similarity (Eq. 15) between the planted and inferred backbones $G^{(b)}_{planted}$ and $G^{(b)}_{MDL}$, using the global and local MDL objectives in Eq.s II.1 and 5 respectively. The difficulty of the reconstruction task is varied by changing the noise parameter $\gamma=\theta_{1}/\theta_{0}$, which is the ratio of the means for the non-backbone and backbone weight distributions, and the degree $k$ of the nodes in the underlying random regular directed graph with $N=100$ nodes. The synthetic networks are drawn from the global Bayesian generative model with geometric weights discussed in Sec. II.3. (b) Reconstruction performance for synthetic networks drawn from the local Bayesian generative model with geometric weights discussed in Sec. II.3. (c) Inverse compression ratio (Eq. 6) for the same set of experiments as in panel (a). (d) Inverse compression ratio for the same experiments as in panel (b). Markers indicate averages over 100 simulations and error bars represent one standard error in the mean.
@@
As a first set of experiments to test our backboning methods, we examine the ability of these methods to reconstruct planted backbone structure in synthetic data simulated from the Bayesian generative models described in Sec. II.3. In the experiments we first generate an unweighted, directed graph with a fixed number of out-edges $k$ for each node $i$. The focus of the MDL backboning methods is on the edge weight distributions, and degree fluctuations simply provide noise in the reconstruction results so degrees are set to be uniform across nodes. We then generate the weights for this network using the global and local generative schemes described in Sec. II.3, with the priors $P(\pi_{b})$ and $P(\theta_{0})$ (for the global generative model) or $P(\theta_{0}(i))$ (for the local generative model) uniform over the range $[0,1]$ and $P(w_{ij}|\theta_{b_{ij}})$ or $P(w_{ij}|\theta_{b_{ij}}(i))$ set to a geometric distribution. As a measure of how “noisy” the backbone is—in other words, how similar the backbone weights and non-backbone weights are—we fix a parameter $\gamma=\theta_{1}/\theta_{0}$ for each simulation which is equivalent to the ratio of the mean non-backbone edge weight and mean backbone edge weight. We therefore set $\theta_{1}=\gamma\theta_{0}$ for each simulation to fix the ratio of the mean weights.
@@
We test the reconstruction and compression capability for the global and local MDL backboning algorithms in Sec.s II.1 and II.2 using two measures. The first is the Jaccard similarity index between the planted backbone $G^{(b)}_{\text{planted}}$ generated from the model and the inferred backbone $G^{(b)}_{\text{MDL}}$ for each method. Since the networks $G$ are represented as edge sets, the Jaccard similarity index between a network $G_{1}$ and a network $G_{2}$ can be computed as
@@
The Jaccard similarity in this case tells us how well the global and local MDL backboning algorithms can recover planted structure from their own (approximate) generative models, and falls in the range $[0,1]$ with 00 indicating an inferred empty or complete graph and $1$ indicating perfect recovery. There are many alternative suitable measures for computing the graph similarity felippe2024network , but in practice they give the same qualitative trends as the Jaccard similarity for this example. The second measure we use to evaluate the methods is the inverse compression ratio $\eta$ (Eq. 6), which tells us how well the methods compress the synthetic network data.
Fig. 2 shows the results of our reconstruction experiments, which were run for random directed regular graphs with $N=100$ nodes, degree $k\in\{5,20,100\}$, and varying noise levels $\gamma\in[0,1]$. Both the global (left column of panels) and local (right column of panels) generative models were used, to determine the extent to which each backboning method is robust under model misspecification. We can see from Fig. 2(a) that, as expected, the reconstruction performance becomes worse as we increase the noise level $\gamma$. At $\gamma\approx 10^{-3}$ we observe near perfect backbone recovery for both methods in the high degree regime, but at $\gamma\approx 1$ we see that neither inferred backbone has any significant shared structure with the planted backbone. We can see that the global method is insensitive to degree $k$ while the local method is sensitive to $k$, since lower $k$ will result in greater weight fluctuations within the node neighborhoods but less significant fluctuations at the global level since $E=500$ even for the lowest degree value ($k=5$) studied. We also see that, as expected, the global MDL method has better performance in the low noise regime than the local MDL method, since it is (approximately) the Bayes-optimal algorithm. However, for higher levels of noise $\gamma$ we see that the local backboning method achieves better performance, which improves for smaller degrees $k$. This is because in this regime the global backboning method tends to infer very sparse backbones, but the local method can still find backbones with a moderate level of connectivity due to considering within-neighborhood weight fluctuations rather than global weight fluctuations.
@@
We plot the inverse compression ratio (Eq. 6) versus the noise level $\gamma$ in Fig. 2(c) and Fig. 2(d) for the global and local generative models respectively. We see that for the global generative model the global MDL method outperforms the local MDL method, and for the local generative model the local MDL method outperforms the global MDL method. This is consistent with the global and local methods being approximately Bayes-optimal for the global and local Bayesian models respectively, as discussed in Sec. II.3. In both cases substantial compression is possible using both methods, and the networks are more compressible for smaller degrees $k$.
The results in Fig. 2 are consistent with the arguments in Sec. II.3 for the correspondence between the proposed MDL methods and Bayesian geometric generative models of network weights. They also demonstrate the capability for these methods to identify meaningful backbone structure and provide considerable compression of network data in the presence of noise. In the next section we compare the MDL methods against existing methods on synthetic data with various levels of planted global and local homogeneity in the edge weights.
@@ III.2 Comparison on Synthetic Networks
Figure 6: Global and local backbones across a real-world network corpus. Global and local MDL backbones were inferred using the objectives of Eq. II.1 and Eq. 5 respectively, for the set of real networks described in Sec. III.3. We plot (a) the fraction of initial edges retained in the local MDL backbone and global MDL backbone and (b) the inverse compression ratio (Eq. 6) for each backbone examined. We only include examples for which both the global and local MDL methods inferred a non-empty backbone, noting the empty backbones in the axis labels. In both panels we plot the line $y=x$ for reference and color the examples for which the global (local) backbone returned a higher value with blue (red). We also indicate with different markers the domain of each example network.
@@
As discussed in the Introduction, the MDL methods we propose in this paper are fully nonparametric, in contrast with existing methods for network backboning which either constrain the topology of the backbone to a specific structure such as a tree, require the specification of the desired number of edges $E^{(b)}$ for the backbone, or require a significance level at which to retain edges under some null model for the weights serrano2009extracting ; foti2011nonparametric ; marcaccioli2019polya ; yassin2023evaluation . This makes direct comparison with existing methods challenging.
@@
We therefore select a few popular existing methods for global and local backboning for comparison, with parameters fixed as in the original studies. The first method we use for comparison is the Disparity Filter serrano2009extracting , which is a principled inference-based method for local backboning that selects edges below a pre-specified significance level $\alpha$ within each node neighborhood for the backbone. The null model used for comparison is one in which the weight assignment is uniform conditioned on the node degree and strength. For more direct comparison with our own local method, for one baseline we fix $E^{(b)}=|G^{(b)}_{\text{local}}|$ as the number of edges to retain in the Disparity Filter and set $\alpha$ to obtain the desired number of edges. For the other Disparity Filter baseline, we set the p-value threshold to the common value of $\alpha=0.05$. We constrain the method to focus on out-neighborhoods for these synthetic examples.
@@
For the other baseline, we compare against a popular global backboning method called the High Salience Skeleton grady2012robust . This method computes the saliency of a link based on its occurrence frequency in the shortest path trees rooted at each node, giving a succinct global view of edge importance. It is observed that the saliency distribution of links in real networks is highly bimodal, so the skeleton is not sensitive to the specific choice of the saliency cutoff for the backbone, which is chosen to be the center of the saliency distribution. The same methodology is applied here for the baseline. Due to its computation of the shortest path tree for each node, the High Salience Skeleton can become too computationally expensive for large networks, so for $N>10000$ we randomly sample the shortest path trees for $10000$ nodes to estimate the link saliency.
@@
The synthetic networks used for the comparison experiments were generated using the following procedure. First, as in Sec. III.1, an unweighted network of $N$ nodes is generated as a $k$-regular random directed graph with the specified out-degree $k$. Given an input parameter $W$ for the total edge weight, all edges are then assigned weight $1$ and the excess weight $W-Nk$ is distributed across the nodes’ out-neighborhoods using a symmetric Dirichlet-Multinomial distribution with concentration parameter $h_{\text{str}}$. This allows us to tune the level of homogeneity in the out-strengths $\{s_{i}\}$ across nodes with a single parameter $h_{\text{str}}$: $h_{\text{str}}\to 0$ places all the excess edge weight in a single out-neighborhood $\partial_{i}$ to give a highly heterogeneous strength distribution, and $h_{\text{str}}\to\infty$ distributes all the weight equally across nodes to give a perfectly homogeneous strength distribution. We then distribute the total excess edge weight $s_{i}-k$ among the out-edges in the neighborhood $\partial_{i}$ using another symmetric Dirichlet-Multinomial distribution, this time with concentration parameter $h_{\text{neig}}$. $h_{\text{neig}}\to 0$ places all the excess edge weight on a single edge $w_{ij}$ to give a highly heterogeneous weight distribution in each neighborhood, and $h_{\text{neig}}\to\infty$ distributes all the weight equally across edges to give a perfectly homogeneous strength distribution. The synthetic networks therefore depend on the four parameters $\{N,W,h_{\text{str}},h_{\text{neig}}\}$, which we vary in our experiments.
@@
We compare backboning methods using multiple indicators, similar to the methodology used in yassin2023evaluation . We compare the fraction of edges $E^{(b)}/E$ and total weight $W^{(b)}/W$ appearing in the backbone to get a sense of how sparse the inferred backbones are with each method. We also compute the inverse compression ratio (Eq. 6) using the global description length (Eq. II.1) for our global MDL method and the High Salience Skeleton, and the local description length (Eq. 5) for our local MDL method and the Disparity Filter variants to see how well each method compresses the data according to the global and local encodings. To examine the extent to which the backbones maintain global connectivity we compute the fraction of nodes $N^{(b)}/N$ with non-zero degree in the backbone and the relative reachability $R$, defined by
@@
In Fig. 3 we plot the fraction of edges (top row) and weight (bottom row) retained in the different backbones of the different backboning methods versus the synthetic model parameters. Unless otherwise specified as an independent variable, in all simulations we set $N=1000$, $W/N=1000$, $k=50$, and $h_{\text{str}}=h_{\text{neig}}=0.1$. There are a few major trends we can observe here. The first is that when $h_{\text{neig}}=0.1$ is held constant at a moderate value (first two columns in Fig. 3), both MDL methods consistently retain a similar number of edges, which is the highest among all methods (an average of around $20\%$ of all edges). We can also see that the High Salience Skeleton consistently retains the fewest edges (an average of around $2\%$ of all edges) in these cases, while the Disparity Filter with an $\alpha=0.05$ significance level finds an intermediate number of backbone edges (an average of around $8\%$ of all edges). We can also see that the local MDL method consistently retains a higher total backbone weight $W^{(b)}$ for the same number of edges $E^{(b)}$ as the Disparity Filter when the latter is set to have the same edge count. The three local backboning methods are highly sensitive to variations in the heterogeneity of weights within the node neighborhoods $h_{\text{neig}}$, while the global MDL method and the High Salience Skeleton are less sensitive to these fluctuations. The local backboning methods all have a sudden drop in the number of edges and total weight they retain at $h_{\text{neig}}\approx 1$, which is the point at which the within-neighborhood variance in the edge weights drops significantly according to this Dirichlet concentration parameter. (This low level of weight homogeneity within neighborhoods leaves little statistical evidence to retain any edges in the backbone.) The local MDL method is the least sensitive to these fluctuations among the local backboning methods.
@@
While the size of the backbones is an important factor for efficient downstream graph computations, it is also important that the sparsified networks retain their global connectivity. In the top row of Fig. 4 we plot the fraction of non-isolated nodes in the backbone (i.e. nodes with degree 00) versus the three model parameters. We can observe that all backbones except the High Salience Skeleton have zero isolated nodes for a large portion of the parameter space in panels (a) and (b), with higher levels of node isolation for very low levels of strength homogeneity $h_{\text{str}}\lesssim 10^{-2}$ due to sparser backbones across the board. In panel (c) we again see strong fluctuations for the local backboning methods—particularly for the Disparity Filter at $\alpha=0.05$—with the MDL method again performing the best out of the local methods and the global MDL method outperforming the High Salience skeleton in terms of node connectivity. The bottom row shows similar connectivity patterns across methods but this time with respect to reachability (Eq. 16). Here the Disparity Filter has a slight edge over the local MDL method in most experiments for the same number of edges and consistently worse reachability for $\alpha=0.05$. The High Salience Skeleton has exactly zero reachabilty in all experiments, suggesting that although this method highlights important edges for global routing it may tend to construct backbones that are too sparse to maintain global connectivity in the network. All methods except the global MDL method perform poorly in the regime of high neighborhood weight homogeneity, becoming too sparse to have a high reachability value.
In Fig. 5 we plot a few additional metrics against the synthetic model parameters. In the top row, we can see that (as expected) the global MDL method is the most compressive among the global methods and the local MDL method is the most compressive among the local methods. Although the local methods tend to be more compressive in general, for high strength homogeneity (panel (b)) we find a slight compressibility advantage for the global method, suggesting that for networks with simple weight distributions one may favor the global MDL method over the local MDL method due to the additional model complexity of the latter. All methods exhibit worse compression as the weights become more homogeneous, since there is little statistical evidence for a separate backbone of high weight edges. In the bottom row we plot the Hellinger distance between the node strength distributions in the full network and the backbone (Eq. 17), which shows similar performance for the global MDL method, local MDL method, and Disparity Filter with $E_{Local}$, while the Disparity Filter at $\alpha=0.05$ has slightly worse performance (e.g. a higher discrepancy in the strength distributions). In panel (f), however, we see that the MDL methods are much more robust to neighborhood weight heterogeneity fluctuations than the other methods, maintaining relatively similar strength distributions in the inferred backbones as the original networks.
@@
In Appendix A we plot the sizes of the inferred backbones versus the number of nodes in the original network (Fig. 8), finding a very weak relationship with this model parameter. We also show the runtime of the different backboning methods versus the network size, in Fig. 9. The slightly superlinear scaling of the MDL runtimes is consistent with the theoretical scaling of $O(E\log E)$ discussed in Sec. II.1 and Sec. II.2. We also observe a substantially higher computational complexity for the High Salience Skeleton due to the computation of the shortest path trees.

### v3

@@ III.1 Synthetic Backbone Reconstruction
Figure 2: Reconstruction of planted backbone structure in synthetic network data. (a) Network Jaccard similarity (Eq. 21) between the planted and inferred backbones $G^{(b)}_{planted}$ and $G^{(b)}_{MDL}$, using the microcanonical global and local MDL objectives in Eq.s 13 and 14 respectively. The difficulty of the reconstruction task is varied by changing the noise parameter $\gamma=\theta_{1}/\theta_{0}$—the ratio of the means for geometric weight distributions in the non-backbone and backbone—as well as the degree $k$ of the nodes in the underlying random regular directed graph with $N=100$ nodes. The synthetic networks in this panel are drawn from the global canonical model in Eq. II.1 with geometric weights. (b) Network Jaccard similarity between the planted and inferred backbones, for networks generated with the local canonical model of Eq. II.1 with geometric weights. (c) Inverse compression ratio (Eq. 20) for the same set of experiments as in panel (a). (d) Inverse compression ratio for the same experiments as in panel (b). Markers indicate averages over 100 simulations and error bars represent one standard error in the mean.
@@
As a first set of experiments to test our backboning methods, we examine the ability of these methods to reconstruct planted backbone structure in synthetic data simulated from the canonical generative models described in Sec. II.1. In the experiments we first generate an unweighted, directed graph with a fixed number of out-edges $k$ for each node $i$. The focus of the MDL backboning methods is on the edge weight distributions, and degree fluctuations simply provide noise in the reconstruction results so degrees are set to be uniform across nodes. We then generate the weights for this network using the global and local canonical generative schemes described in Sec. II.1, with the priors $P(\pi_{b})$ and $P(\theta_{0})$ (for the global generative model) or $P(\theta_{0}(i))$ (for the local generative model) uniform over the range $[0,1]$ and $P(w_{e}|\theta_{b_{e}})$ or $P(w_{e}|\theta_{b_{e}}(i))$ set to a geometric distribution. As a measure of how “noisy” the backbone is—in other words, how similar the backbone weights and non-backbone weights are—we fix a parameter $\gamma=\theta_{1}/\theta_{0}$ for each simulation which is equivalent to the ratio of the mean non-backbone edge weight and mean backbone edge weight. We set $\theta_{1}=\gamma\theta_{0}$ for each simulation to fix the ratio of the mean weight, with the exact mean weights $1/\theta_{1}$ and $1/\theta_{0}$ (and consequently the weight variances) varying for each trial based on the sampled value $\theta_{0}\sim\text{Uniform}([0,1])$.
@@
We test the reconstruction and compression capability for the global and local MDL backboning methods in Sec. II.2 using two measures. The first is the Jaccard similarity index between the planted backbone $G^{(b)}_{\text{planted}}$ generated from the model and the inferred backbone $G^{(b)}_{\text{MDL}}$ for each method. Since the networks $G$ are represented as edge sets, the Jaccard similarity index between a network $G_{1}$ and a network $G_{2}$ can be computed as
@@
The Jaccard similarity in this case tells us how well the global and local MDL backboning algorithms can recover planted structure from their own (approximate canonical) generative models, and falls in the range $[0,1]$ with 00 indicating an inferred empty or complete graph and $1$ indicating perfect recovery. There are many alternative suitable measures for computing the graph similarity such as the network mutual information felippe2024network , but in practice they give the same qualitative trends as the Jaccard similarity for this example. The second measure we use to evaluate the methods is the inverse compression ratio $\eta$ (Eq. 20), which tells us how well the methods compress the synthetic network data.
Fig. 2 shows the results of our reconstruction experiments, which were run for random directed regular graphs with $N=100$ nodes, degree $k\in\{5,20,100\}$—allowing self-loops—and varying noise levels $\gamma\in[0,1]$. Both the global (left column of panels) and local (right column of panels) generative models were used, to determine the extent to which each backboning method is robust under model mis-specification. We can see from Fig. 2(a) that, as expected, the reconstruction performance becomes worse as we increase the noise level $\gamma$. At $\gamma\approx 10^{-3}$ we observe near perfect backbone recovery for both methods in the high degree regime, but at $\gamma\approx 1$ we see that neither inferred backbone has any significant shared structure with the planted backbone. We can see that the global method is insensitive to degree $k$ while the local method is sensitive to $k$, since lower $k$ will result in greater weight fluctuations within the node neighborhoods but less significant fluctuations at the global level since $E=500$ even for the lowest degree value ($k=5$) studied. We also see that, as expected, the global MDL method has better performance in the low noise regime than the local MDL method, since it is (approximately) the Bayes-optimal algorithm. However, for higher levels of noise $\gamma$ we see that the local backboning method achieves better performance, which improves for smaller degrees $k$. This is because in this regime the global backboning method tends to infer very sparse backbones, but the local method can still find backbones with a moderate level of connectivity due to considering within-neighborhood weight fluctuations rather than global weight fluctuations.
@@
We plot the inverse compression ratio (Eq. 20) versus the noise level $\gamma$ in Fig. 2(c) and Fig. 2(d) for the global and local generative models respectively. We see that for the global generative model the global MDL method outperforms the local MDL method, and for the local generative model the local MDL method outperforms the global MDL method. This is consistent with the global and local methods being approximately Bayes-optimal for the global and local generative models respectively, due to the asymptotic equivalence discussed in Appendix A. In both cases substantial compression is possible using both methods.
In Appendix B, we demonstrate numerically the optimality of the greedy algorithm in similar reconstruction tests on much smaller networks where exact enumeration over backbones is possible, to support the calculations in Sec. II.3.
The results in Fig. 2 demonstrate the capability for these methods to identify meaningful backbone structure and provide considerable compression of network data in the presence of noise and model mis-specification. In the next section we compare the MDL methods against existing methods on synthetic data with various levels of planted global and local homogeneity in the edge weights.
@@ III.2 Comparison on Synthetic Networks
Figure 3: Six backboning methods applied to a network of workplace contacts. All six methods used for the tests in Sec. III were applied to the workplace contacts network of genois2015data , obtained from the Netzschleuder repository netz along with the other networks described in Sec. III.3. Edges in the backbone are highlighted in color, while all other edges are black. Edges are scaled proportionally to weight. The fraction of edges $E^{(b)}/E$ and nodes $N^{(b)}/N$ retained in the backbone are listed alongside each backboning method above the corresponding plot. The backbone densities and node connectivity (in terms of number of edges retained and number of non-isolated nodes in the backbone) observed for this network are fairly consistent with those observed for other real networks. The percolation backbone has the highest connectivity at the cost of a higher density; the disparity filter with $\alpha=0.05$ and the high salience skeleton have the lowest density at the cost of the lowest connectivity; and the three methods of the top row (global MDL, local MDL, and disparity filter with $E$ equal to the local MDL method) retain most of the node connectivity with lower edge densities.
@@
Figure 6: Global and local backbones across a real-world network corpus. Global and local MDL backbones were inferred using the objectives of Eq. 13 and Eq. 14 respectively, for the set of real networks described in Sec. III.3. We plot (a) the fraction of initial edges retained in the local MDL backbone and global MDL backbone and (b) the inverse compression ratio (Eq. 20) for each backbone examined. We only include examples for which both the global and local MDL methods inferred a non-empty backbone, noting the empty backbones in the axis labels. In both panels we plot the line $y=x$ for reference and color the examples for which the global (local) backbone returned a higher value with blue (red). We also indicate with different markers the domain of each example network.
@@
The MDL methods we propose in this paper are fully nonparametric, in contrast to many existing methods which require the specification of the desired number of edges $E^{(b)}$ for the backbone or require a significance level at which to retain edges under some null model for the weights serrano2009extracting ; foti2011nonparametric ; marcaccioli2019polya ; yassin2023evaluation . This makes direct comparison with existing methods challenging, since it is unclear how to choose free parameters for methods that require them.
@@
We therefore select a few popular existing methods for global and local backboning for comparison, with parameters fixed to common values. The first method we use for comparison is the Disparity Filter serrano2009extracting , which is a principled inference-based method for local backboning that selects edges below a pre-specified significance level $\alpha$ within each node neighborhood for the backbone. The null model used in the Disparity Filter is one in which the weight assignment is uniform conditioned on the node degree and strength. For more direct comparison with our own local method, for one baseline we fix $E^{(b)}=|G^{(b)}_{\text{local}}|$ as the number of edges to retain in the Disparity Filter and set $\alpha$ to obtain the desired number of edges. For the other Disparity Filter baseline, we set the p-value threshold to the common value of $\alpha=0.05$. We constrain the method to focus on out-neighborhoods for these synthetic examples.
@@
For the first global backboning baseline, we use the High Salience Skeleton grady2012robust , which computes the saliency of a link based on its occurrence frequency in the shortest path trees rooted at each node, giving a succinct global view of edge importance. It is observed that the saliency distribution of links in real networks is highly bimodal, so the skeleton is not sensitive to the specific choice of the saliency cutoff for the backbone, which is chosen to be the center of the saliency distribution ($0.5$). The same methodology is applied here for the baseline. Due to its computation of the shortest path tree for each node, the High Salience Skeleton can become too computationally expensive for large networks, so for $N>10000$ we randomly sample the shortest path trees for $10000$ nodes to estimate the link saliency. For the other global backboning baseline, we use the weighted percolation threshold method li2007transport ; haimovici2013brain , which consists of adding the edges to the network in decreasing order of weight until the network forms a fully connected component. In Fig. 3 we plot the results of all six measures applied to the workpace contacts dataset of genois2015data , obtained using the procedures described in Sec. III.3.
@@
The synthetic networks used for the comparison experiments were generated using the following procedure. First, as in Sec. III.1, an unweighted network of $N=1000$ nodes is generated as a $k$-regular random directed graph with the specified out-degree $k=50$. Given an input parameter $W$ for the total edge weight, all edges are then assigned weight $1$ and the excess weight $W-Nk$ is distributed across the nodes’ out-neighborhoods using a symmetric Dirichlet-Multinomial distribution with concentration parameter $h_{\text{str}}$. This allows us to tune the level of homogeneity in the out-strengths $\{s_{i}\}$ across nodes with a single parameter $h_{\text{str}}$: $h_{\text{str}}\to 0$ places all the excess edge weight in a single out-neighborhood $\partial_{i}$ to give a highly heterogeneous strength distribution, and $h_{\text{str}}\to\infty$ distributes all the weight equally across nodes to give a perfectly homogeneous strength distribution. We then distribute the total excess edge weight $s_{i}-k$ among the out-edges in the neighborhood $\partial_{i}$ using another symmetric Dirichlet-Multinomial distribution, this time with concentration parameter $h_{\text{neig}}$. $h_{\text{neig}}\to 0$ places all the excess edge weight on a single edge $w_{ij}$ to give a highly heterogeneous weight distribution in each neighborhood, and $h_{\text{neig}}\to\infty$ distributes all the weight equally across edges to give a perfectly homogeneous weight distribution within the neighborhoods. The synthetic networks therefore depend on the four parameters $\{N,W,h_{\text{str}},h_{\text{neig}}\}$, which we vary in our experiments.
@@
We compare backboning methods using multiple indicators, similar to the methodology used in yassin2023evaluation . We compare the fraction of edges $E^{(b)}/E$ and total weight $W^{(b)}/W$ appearing in the backbone to get a sense of how sparse the inferred backbones are with each method. We also compute the inverse compression ratio (Eq. 20), using the global description length (Eq. 13) for the global methods (global MDL, High Salience Skeleton, Percolation) and the local description length (Eq. 14) for the local methods (local MDL, Disparity Filters) to see how well each method compresses the data. To examine the extent to which the backbones maintain global connectivity we compute the fraction of nodes $N^{(b)}/N$ with non-zero degree in the backbone and the relative reachability $R$, defined by
@@
In Fig. 4 we plot the fraction of edges (top row) and weight (bottom row) retained in the different backbones of the different backboning methods versus the synthetic model parameters. Unless otherwise specified as an independent variable, in all simulations we set $N=1000$, $W/N=1000$, $k=50$, and $h_{\text{str}}=h_{\text{neig}}=0.1$. There are a few major trends we can observe here. The first is that when $h_{\text{neig}}=0.1$ is held constant at a moderate value (first two columns in Fig. 4), both MDL methods consistently retain a similar number of edges (an average of around $20\%$ of all edges), which is the highest among all methods for $W/N\gtrsim 10^{3}$. We can also see that the High Salience Skeleton consistently retains the fewest edges (an average of around $2\%$ of all edges) in these cases, while the Disparity Filter with an $\alpha=0.05$ significance level and the percolation threshold find an intermediate number of backbone edges (an average of around $8\%$ and $12\%$ of all edges respectively). We also observe that the local MDL method consistently retains a higher total backbone weight $W^{(b)}$ for the same number of edges $E^{(b)}$ as the Disparity Filter when the latter is set to have the same edge count. The three local backboning methods are highly sensitive to variations in the heterogeneity of weights within the node neighborhoods $h_{\text{neig}}$ (panels (c) and (f)), while the global MDL method and the High Salience Skeleton are less sensitive to these fluctuations and the percolation method is highly insensitive to the fluctuations. The local backboning methods all have a sudden drop in the number of edges and total weight they retain at $h_{\text{neig}}\approx 1$, which is the point at which the within-neighborhood variance in the edge weights drops significantly according to this Dirichlet concentration parameter. This high level of weight homogeneity within neighborhoods leaves little statistical evidence to retain any edges in the backbone, since no subset of weights are significantly higher than the rest.
@@
While the size of the backbones is an important factor for efficient downstream graph computations, it is also important that the sparsified networks retain their global connectivity. In the top row of Fig. 5 we plot the fraction of non-isolated nodes in the backbone (i.e. nodes with degree 00) versus the three model parameters. We can observe that all backbones except the High Salience Skeleton have nearly zero isolated nodes for a large portion of the parameter space in panels (a) and (b), with higher levels of node isolation for very low levels of strength homogeneity $h_{\text{str}}\lesssim 10^{-2}$ due to sparser backbones across the board. In panel (c) we again see strong fluctuations for the local backboning methods—particularly for the Disparity Filter at $\alpha=0.05$—with the MDL method again performing the best out of the local methods and the global MDL method outperforming the High Salience skeleton in terms of node connectivity. The percolation-based method, by definition, has zero isolated nodes for all parameter settings, at the cost of consistently high fractions of edges being retained in this homogeneous weight regime (Fig. 4c).
The bottom row of Fig. 5 shows similar connectivity patterns across methods but this time with respect to reachability (Eq. 22). Here the Disparity Filter has a slight edge over the local MDL method in most experiments for the same number of edges and consistently worse reachability for $\alpha=0.05$. The High Salience Skeleton has exactly zero reachabilty in all experiments, suggesting that although this method highlights important edges for global routing it may tend to construct backbones that are too sparse to maintain global connectivity in the network. The percolation backbone, despite enforcing a single weakly connected component, does not fully retain directed connectivity as required by $R(G^{(b)})$, scoring lower than other methods in most tests. All methods except the global MDL and percolation methods have poor connectivity in the regime of high neighborhood weight homogeneity, becoming too sparse to have a high reachability value when there is little signal to distinguish the neighborhood backbones.
@@
In Fig. 10 in Appendix C, we plot a few additional metrics against the synthetic model parameters, to understand how compressive each method is and how well each preserves the strength distribution of the original network. We also plot the sizes of the inferred backbones versus the number of nodes in the original network in the same Appendix (Fig. 11), finding a very weak relationship with this model parameter. Finally, we show the runtime of some different backboning methods versus the network size, in Fig. 12. The slightly superlinear scaling of the MDL runtimes is consistent with the theoretical scaling of $O(E\log E)$ discussed in Sec. II.3. We also observe a substantially higher computational complexity for the High Salience Skeleton due to the computation of the shortest path trees.